# Learning Program Behavioral Models from Synthesized Input-Output Pairs

### v1

@@ 4.5. Results
Domain-specific tokenizers (as provided by output grammars) are better than generic tokenizers.
@@ 5. Limitations
Modelizer, will not be able to create models for all sorts of programs. The complexity of the whole program logic or its individual capabilities can be too high to be modeled by the sequence-to-sequence models. The transformer models themselves have a quadratic computational complexity. For the input sequence length $n$, hidden dimension size $h$, and number of layers in the neural network $L$, we get the following computational complexity:
Time $\textit{Time Complexity}:O\bigl{(}L\times(n^{2}d+nd^{2})\bigr{)}$
Space $\textit{Space Complexity}:O\bigl{(}L\times(n^{2}+nd^{2})\bigr{)}$
This means that modeling of extremely long input-output will be challenging and impractical. The models that are learned by Modelizer are not intended to replace programs. We see Modelizer as a tool that can help to understand the behavior of the program and assist in the program testing and debugging.
@@
Similar to many Transformer-based systems, Modelizer is sensitive to the length of the processed sequences. One of the factors that impact translation accuracy is the synthesis of training sequences. The grammar-based fuzzers used in the implementation of our Generators, by default, tend to produce short sequences. To overcome this problem we have introduced the sliding non-terminal expansion window, which partially resolves this issue. While the total token quantity per input and output sequences reached values similar to real-world samples, it does not guarantee a correlation for the input/output feature distributions between synthesized and real-world samples. At the same time, such a distribution can be learned from existing input samples (Soremekun et al., 2022). This approach introduces three requirements: such inputs have to be (1) available before input generation, (2) parsable with selected grammar, and (3) the correct values have to be set for the minimal and maximal number of non-terminal expansions in the generator. In the current study, we intentionally have not followed this approach to check Modelizer’s capabilities in learning behavior models of arbitrary systems given only the formal specification of such systems.
@@
Certain subjects, specifically the output data types like KQL, required the implementation of custom parsers due to the unavailability of third-party parsers at the time this publication was prepared. The implementation of such parsers is a time-consuming and error-prone task. The third-party pre-trained tokenizers suboptimally partition the corresponding sequences, by unnecessarily splitting atomic tokens into sub-tokens similar to the constructions that are found in natural language sequences, misinterpreting certain characters like punctuation symbols, which cover part of the format syntax. Partially, such an issue can be mitigated by defining custom post-processing routines that will refine the outcome of the tokenization phase, but, such a list of regular expressions and string replacement rules can quickly become very long, create dependencies and conflicts in the order or application, what makes them hard to maintain. Such measures do not simplify the tokenization process but rather move the complexity from the tokenization phase to the post-processing phase. While the custom subject-specific parser-tokenizers are robust and produce high-quality tokens, we do not think that automated model learning problems can benefit from this approach in the long run, specifically for cases when complex system behavior has to be modelized by chaining the behavior models of sub-systems. We think that such a domain-specific tokenizer can be automatically created by wrapping existing tokenization strategies like byte-pair encoding, word-piece, or sentence-piece tokenization with a reinforcement learning loop. Alternatively, we are looking forward to solutions that would be able to mine the input-output grammars from the PUTs and automatically infer the parser-tokenizers from them.
@@
Given the correct grammar definition, our generators which rely on grammar-based fuzzers, are very efficient in producing syntactically correct inputs. However, the PUT still can reject these inputs because grammar-based fuzzers cannot guarantee that the generated input will be semantically valid and pass all parsing guards in the PUT. Thus, an additional refinement of the synthesized data might be required. Steinhöfel et al. (Dominic Steinhöfel and Andreas Zeller, 2022) have proposed a technique and a formal language that allows manual specification of the semantic constraints inferred from the system specification, domain knowledge, and human expertise. Their approach, which builds on a modern Z3 SMT solver, can automatically refine the semantically invalid inputs using encoded constraints. However, fuzzing with semantic constraints is a computationally expensive task, and it slows down the training dataset formation on a large scale. One type of inputs that we synthesized, and which required such a refinement, were SQL queries with a JOIN condition. Using our SQL grammar, the GrammarFuzzer was able to produce SQL queries similar to the one given below:
@@
Once PlaceholderProcessor updates all placeholders with unique identifiers the query starts to look like this:
@@
The refined query is semantically invalid. To overcome this issue, we implemented an additional string refinement step, which parses the synthesized input and corrects the identifiers. While such an additional refinement step requires a bit of manual effort, for simple cases like this, it adds an insignificant performance penalty to the input generation process. Thus, the current implementation of the Modelizer framework does not support fuzzing with constraint solving, but we plan to support this functionality in future releases.
@@ 6. Related Work
The evolution of language models and their integration into existing projects and solutions have opened new horizons in automating routine tasks and processes (Chen et al., 2021; Ziegler et al., 2022). The so-called general-purpose large language models (LLMs) like Llama 2 (Touvron et al., 2023) or T5 (Raffel et al., 2020) have been adapted to process code-related tasks by fine-tuning on a huge corpus of publicly available source codes and documentation (Code Llama (Rozière et al., 2024) or CodeT5 (Wang et al., 2021)). Such optimized code-specific models can be further fine-tuned on code generation and understanding downstream tasks. Below we discuss the most relevant applications of neural models in the context of program behavior modeling and the challenges they face:
Behavior Mocking.:
@@
There have already been previous attempts to perform behavior mocking by training sequence-to-sequence neural networks. TransCoder (Roziere et al., 2020) represents two sequence-to-sequence models for cross-language translations of code written in C++, Java, and Python with additional support for back translations. Armengol-Estape et al. (Armengol-Estapé and O’Boyle, 2022) train transformer models to replicate the behavior of the GCC compiler. The training data is obtained by selecting C functions from the Anghabench benchmark and compiling them to x86 assembly code. The learned models can correctly compile the C code in less than half of the cases, making replacing compilers with neural machine Translation models invisible in practice. Zaremba et al., in their “Learning To Execute” publication (Zaremba and Sutskever, 2015), try to predict the execution results of short Python programs performing simple arithmetic operations and echo printing by training an LSTM (Hochreiter and Schmidhuber, 1997) model of a Python interpreter on a set of synthesized programs. These works have shown that the current generation of neural network architectures cannot correctly learn the full behavior of such complex programs as compilers or interpreters. While learned models cannot fully replace complex programs, our approach has demonstrated promising results in modeling linear or linearithmic programs. In contrast to previous approaches that rely on sequence-to-sequence models to achieve behavior mocking, Shen. et al and their Konure (Shen and Rinard, 2021) system relies on the active learning approach to mine the behavior model of applications that interact with relational databases. Konure synthesizes inputs with database content simultaneously, monitors the traffic exchange between the database and the application, and inspects the results of the database queries to continuously refine the application behavior hypothesis. Using the refined application model, Konure can generate the Python code for database querying that is semantically equivalent to the original application code, which might be written in a different programming language. Similar to Konure, our approach learns the program models from the input-output pairs. While Konure is designed to model the database interaction behavior of applications, our approach does not have such a limitation and can be used to either model the behavior of the whole system or its components and individual functionality.
Reversing Computation.:

### v2

@@ 4.5. Results
Structure-aware tokenizers are more suitable for behavior learning than pre-trained or unsupervised tokenizers.
RQ6: Could LLMs pre-trained on code mock programs?
In this research question, we examined the ability of LLMs to reproduce program behavior and compared their performance with custom domain-specific models that we have previously trained. At first, we decided to perform zero-shot prompting of LLMs pre-trained on code. However, selected LLMs, CodeLlama (Rozière et al., 2024) and CodeGemma (Team et al., 2024), were instruction fine-tuned by their vendors, so we prepended every prompt with the following instruction template:
You are a converter program that translates {SOURCE} to {TARGET}. Translate the following data: {DATA}
Due to the zero-shot prompting strategy, the instruction does not contain any task-specific examples but, instead, instruments the model with contextual knowledge. At the time of querying, {SOURCE}, {TARGET}, and {DATA} placeholders were replaced with actual values. We summarized the zero-shot prompting results in Table 10. We recorded zero or close to zero Exact Match and Close Match scores for all evaluated subjects. It means that selected LLMs were unable to replicate neither actual program behavior nor the behavior of the hypothetical inverse program. From BLEU score values, we can infer that the CodeGemma model has a limited understanding of Markdown to HTML and HTML to Markdown conversion tasks, while CodeLlama models struggle to provide any valuable results for any data pair and conversion direction combinations without additional fine-tuning. We provide an example of interaction with selected LLMs, including the model outputs, in Table 9. Despite the ability to better understand code-specific tasks due to additional pre-training, these models can not act as program stubs without additional fine-tuning.
Table 9. An example of zero-shot prompting of LLMs pre-trained on code converting Markdown to HTML
Model Input Expected Output CodeLlama Output CodeGemma Output
You are a converter program
that translates Markdown to HTML.
Translate the following data:
## BarkProcessor
<h2>BarkProcessor</h2>
### BarkProcessor.cs
“‘csharp
using System;
using System.Collections.Generic;
<h2><strong><em>
BarkProcessor
</em></strong></h2>
Table 10. Evaluation results for models mimicking Pandoc, msticpy, latexify-py, and py-asciimath via LLMs pre-trained on code.
Program Source Target
Model
Type
Model BLEU
BLEU
Error
NIST WER WIL
Exact
Match
Close
Match
CodeLlama 0.1070 0.0010 1.8602 174.23 97.23 0 0
Markdown HTML Forward CodeGemma 0.4366 0.0029 7.4377 68.24 68.82 0.61 2.17
CodeLlama 0.0875 0.0009 1.4325 266.98 97.02 0 0
Pandoc HTML Markdown Backward CodeGemma 0.5348 0.0028 9.1093 49.84 49.67 0 5.70
CodeLlama 0.0542 0.0009 1.0152 662.57 93.47 0 0
SQL KQL Forward CodeGemma 0.1546 0.0039 2.5782 250.21 84.32 0 0
CodeLlama 0.0338 0.0010 0.6302 644.66 96.77 0 0
msticpy KQL SQL Backward CodeGemma 0.2960 0.0043 5.0848 67.94 52.17 0 0
CodeLlama 0.0767 0.0011 1.4521 266.39 94.65 0 0
PyExpres. Forward CodeGemma 0.1855 0.0032 3.4188 108.77 83.83 0 0
CodeLlama 0.0545 0.0012 0.9835 387.52 94.74 0 0
latexify-py PyExpres. Backward CodeGemma 0.2428 0.0048 4.1169 100.26 76.04 0 0
CodeLlama 0.0076 0.0003 0.3041 988.32 98.10 0 0
AsciiMath Forward CodeGemma 0.0737 0.0047 2.3622 121.12 85.26 0.70 1.20
CodeLlama 0.0077 0.0004 0.3127 894.36 98.37 0 0
AsciiMath Backward CodeGemma 0.1231 0.0045 2.9856 117.92 82.96 0.10 0.10
CodeLlama 0.0023 0.0002 0.2507 1026.46 98.62 0 0
MathML Forward CodeGemma 0.0853 0.0030 2.5428 109.35 81.96 0 0
CodeLlama 0.0037 0.0002 0.5105 155.67 99.34 0 0
MathML Backward CodeGemma 0.0172 0.0006 0.9931 80.27 85.70 0 0
CodeLlama 0.0078 0.0005 0.5531 154.20 99.26 0 0
AsciiMath MathML Hypothet. Forward CodeGemma 0.0190 0.0006 1.0907 78.96 86.87 0 0
CodeLlama 0.0023 0.0002 0.2408 1094.34 98.55 0 0
py-asciimath MathML AsciiMath Hypothet. Backward CodeGemma 0.0412 0.0030 1.5875 140.33 88.26 0 0.50
Blue cells = CodeLlama 7B; purple cells = CodeGemma 7B.
Table 11. Evaluation results for models mocking Pandoc, msticpy, latexify-py, and py-asciimath behaviors queried after fine-tuning.
Program Source Target
Model
Type
Model BLEU
BLEU
Error
NIST WER WIL
Exact
Match
Close
Match
CodeLlama 0.4248 0.0033 6.4448 87.7 71.89 6.78 12.16
Markdown HTML Forward CodeGemma 0.3463 0.0027 5.6852 82.44 76.61 0 0
CodeLlama 0.4148 0.0031 6.4789 95.00 64.99 0 13.05
Pandoc HTML Markdown Backward CodeGemma 0.4701 0.0022 7.9473 51.40 54.69 0 0
CodeLlama 0.4240 0.0071 6.6428 82.81 56.16 0 3.00
SQL KQL Forward CodeGemma 0.7035 0.0045 10.3251 22.59 25.89 0 0
CodeLlama 0.7787 0.0055 11.0942 14.42 21.95 0 15.20
msticpy KQL SQL Backward CodeGemma 0.0896 0.0096 1.2981 427.72 91.98 0 0
CodeLlama 0.4270 0.0054 6.1379 59.58 52.22 0 0
PyExpres. Forward CodeGemma 0.5250 0.0029 7.7616 27.17 39.37 0 0
CodeLlama 0.4777 0.0103 6.7309 82.56 52.97 10.10 14.90
latexify-py PyExpres. Backward CodeGemma 0.8204 0.0048 11.3188 16.55 18.91 0 0
CodeLlama 0.5603 0.0085 8.3336 43.06 41.39 1.40 21.20
AsciiMath Forward CodeGemma 0.5500 0.0080 8.2532 37.55 39.34 0 0
CodeLlama 0.2036 0.0084 3.6731 133.84 76.99 2.00 2.60
AsciiMath Backward CodeGemma 0.3616 0.0076 6.0105 64.78 61.27 0 0
CodeLlama 0.2441 0.0066 4.9843 76.21 69.62 0.50 0.70
MathML Forward CodeGemma 0.2991 0.0072 5.4391 65.28 64.26 0 0
CodeLlama 0.5168 0.0050 4.0192 75.85 69.88 0 0
MathML Backward CodeGemma 0.5542 0.0043 4.4655 63.05 62.49 0 0
CodeLlama 0.5176 0.0049 3.9829 77.77 69.13 0 0
AsciiMath MathML Hypothet. Forward CodeGemma 0.5210 0.0043 4.0966 68.45 64.73 0 0
CodeLlama 0.2892 0.0071 5.3566 68.68 62.83 0.30 3.30
py-asciimath MathML AsciiMath Hypothet. Backward CodeGemma 0.3590 0.0073 6.1543 52.99 57.63 0 0
Outliers
Pandoc Markdown HTML Forward CodeLlama 0.5645 0.0026 4.9441 73.09 60.17 15.88 71.97
Pandoc HTML Markdown Backward CodeLlama 0.2424 0.0028 2.3148 188.1 78.01 0 40.18
latexify-py PyExpres. Backward CodeLlama 0.5306 0.0100 5.065 73.24 46.79 15.4 20.2
py-asciimath AsciiMath Forward CodeLlama 0.8283 0.0058 7.0132 15.54 18.09 1.9 30.4
Blue cells = CodeLlama 7B model; purple cells = CodeGemma 7B model;
yellow cells = CodeLlama 7B with the content part of model input being masked with placeholders.
Table 12. Comparison of fine-tuning times (in minutes) for CodeLlama 7B and CodeGemma 7B across multiple subjects.
Subject CodeGemma CodeLlama Custom 10k Custom 250k Custom 1M
Markdown - HTML 304.6 296.0 26.5 407.4 1359.1
HTML - Markdown 305.7 297.6 27.0 407.2 1354.3
SQL - KQL 517.9 506.8 18.1 451.9 1827.4
KQL - SQL 520.2 508.2 13.4 330.8 1322.3
PyExpres. - LaTeX 1248.8 1278.6 23.8 619.9 2534.4
LaTeX - PyExpres. 1244.4 1281.2 14.9 375.3 1488.2
LaTeX - AsciiMath 113.0 119.9 25.2 642.2 2626.6
AsciiMath - LaTeX 112.8 120.2 13.1 315.8 1318.7
MathML - LaTeX 261.6 256.3 28.0 704.6 2904.5
LaTeX - MathML 263.2 258.5 23.1 570.5 2314.2
AsciiMath - MathML 258.9 253.8 18.1 446.8 1823.4
MathML - AsciiMath 259.1 254.6 22.9 579.5 2374.0
After zero-shot prompting, we independently fine-tuned foundational models for conversion tasks based on every subject-to-conversion direction combination using 10,000 synthesized program input-output pairs to exclude the possibility of cross-subject interference and simplify the comparison with custom subject-specific models. We also reused the same instruction template from zero-shot prompting experiments during the model fine-tuning and inference phases. We recorded significantly higher BLEU score values for all subjects, meaning that models now better understand the mapping between program input-output features, thus forming better syntactically and semantically structured predictions. However, Exact Match scores reflect that pre-trained LLMs cannot serve as in-place mocks for programs. Fine-tuning does not completely prevent models from injecting additional tokens, which is intolerable for programs expecting well-structured inputs. We assume that iteratively refining incorrect model predictions using additional prompting rounds and fine-tuning on more samples could help fix hallucinations of LLMs in behavior-mocking and reverse engineering tasks. Once we additionally masked the content part of the model input with corresponding placeholders in the same way as we did in custom-trained models, we achieved better model performance in half of the subjects with CodeLlama models. These cases are marked as outliers in Table 11.
We also compared the LLM fine-tuning time with training custom models from scratch on different training sets consisting of 10,000 (small), 250,000 (medium), and 1,000,000 (large) entries. While custom model training was performed on a widely used Nvidia RTX 4090 high-end gaming level card, fine-tuning of quantized LLMs required a more powerful data-center grade Nvidia A100 Tensor Core GPU. Despite the dissimilarity in the computation capabilities of selected hardware accelerators, the results in Table 12 indicate that LLM fine-tuning on a small dataset requires time similar to training custom models from scratch using a medium-size dataset with 250,000 entries.
In conclusion, despite our early results in the field of automated behavior extraction, we think that custom-trained models are more suitable for program behavior mocking and reverse engineering of evaluated subjects. We want to point out several core differences. Using LLMs sometimes requires processing data by third parties, which could be undesired or disallowed by terms and conditions, for example, because vendors host models themselves and only provide API access to end-users. Typically, proprietary models require significant computation resources, which are inaccessible to individual developers and small-sized companies. Depending on usage patterns, the paywall could restrict API access to proprietary models and become costly for long-term use. At the same time, our experiments with locally hosted, relatively small by current industry standards LLMs containing 7 billion parameters have demonstrated that such LLMs cannot replace programs in typical usage scenarios. Even after fine-tuning, LLMs still struggle to replicate the behavior identical to the actual program. In contrast, program-specific models learned by Modelizer
  • achieve higher forward and inverse mocking accuracy;
  • have significantly lower computation requirements; and
  • allow for integration into existing deployment or testing scenarios.
at a smaller cost. We foresee a big potential for the co-integration of lightweight program-specific models with large language models and large reasoning models.
Behavior mocking with LLMs is less efficient than training custom models because they are not task-specific, require more resources, and offer similar or lower performance.
@@ 5. Limitations
Modelizer can not create models for all sorts of programs. The program logic or its individual capabilities can be too complex to be modeled by sequence-to-sequence models. The transformer models themselves have a quadratic computational complexity. For the input sequence length $n$, hidden dimension size $h$, and number of layers in the neural network $L$, we get the following computational complexity:
[…]
This means that modeling of extremely long input-output will be challenging and impractical. The models that are learned by Modelizer are not intended to ubiquitously replace programs. We see Modelizer as a tool that can help to understand the behavior of programs and assist in program testing and debugging.
@@
Similar to many Transformer-based systems, Modelizer is sensitive to the length of the processed sequences. One of the factors that impact translation accuracy is the synthesis of training sequences. By default, the grammar-based fuzzers used in the implementation of our Generators tend to produce short sequences. To overcome this problem, we have introduced the sliding non-terminal expansion window, which partially resolves this issue. While the total token quantity per input and output sequences reached values similar to real-world samples, it does not guarantee a correlation for the input/output feature distributions between synthesized and real-world samples. At the same time, such a distribution can be learned from existing input samples (Soremekun et al., 2022). This approach introduces three requirements: such inputs have to be (1) available before input generation, (2) parsable with selected grammar, and (3) the correct values have to be set for the minimal and maximal number of non-terminal expansions in the generator. In the current study, we intentionally have not followed this approach to check Modelizer’s capabilities in learning behavior models of arbitrary systems given only the formal specification of such systems.
@@
Certain subjects, specifically the output data types like KQL, required the implementation of custom parsers due to the unavailability of third-party parsers at the time this publication was prepared. The implementation of such parsers is a time-consuming and error-prone task. The third-party pre-trained tokenizers suboptimally partition the corresponding sequences by unnecessarily splitting atomic tokens into sub-tokens similar to the constructions that are found in natural language sequences, misinterpreting certain characters like punctuation symbols, which cover part of the format syntax. Partially, such an issue can be mitigated by defining custom post-processing routines that will refine the outcome of the tokenization phase, but, such a list of regular expressions and string replacement rules can quickly become very long, create dependencies and conflicts in the order or application, what makes them hard to maintain. Such measures do not simplify the tokenization process but rather move the complexity from the tokenization phase to the post-processing phase. While the custom subject-specific parser-tokenizers are robust and produce high-quality tokens, we do not think that automated model learning problems can benefit from this approach in the long run, specifically for cases when complex system behavior has to be modelized by chaining the behavior models of sub-systems. We think that such a domain-specific tokenizer can be automatically created by wrapping existing tokenization strategies like byte-pair encoding, word-piece, or sentence-piece tokenization with a reinforcement learning loop. Alternatively, we are looking forward to solutions that would be able to mine the input-output grammars from the PUTs and automatically infer the parser-tokenizers from them.
@@
Given the correct grammar definition, our generators, which rely on grammar-based fuzzers, are very efficient in producing syntactically correct inputs. However, the PUT still can reject these inputs because grammar-based fuzzers cannot guarantee that the generated input will be semantically valid and pass all parsing guards in the PUT. Thus, an additional refinement of the synthesized data might be required. Steinhöfel et al. (Steinhöfel and Zeller, 2022) have proposed a technique and a formal language that allows manual specification of the semantic constraints inferred from the system specification, domain knowledge, and human expertise. Their approach, which builds on a modern Z3 SMT solver, can automatically refine the semantically invalid inputs using encoded constraints. However, fuzzing with semantic constraints is a computationally expensive task, and it slows down the formation of training datasets on a large scale. One type of inputs that we synthesized, and which required such a refinement, were SQL queries with a JOIN condition. Using our SQL grammar, the GrammarFuzzer was able to produce SQL queries similar to the one given below:
@@
Once PlaceholderProcessor updates all placeholders with unique identifiers, the query starts to look like this:
@@
The refined query is semantically invalid. To overcome this issue, we implemented an additional string refinement step, which parses the synthesized input and corrects the identifiers. While such an additional refinement step requires a bit of manual effort, it adds an insignificant performance penalty to the input generation process for simple cases like this. Thus, the current implementation of the Modelizer framework does not support fuzzing with constraint solving, but we plan to support this functionality in future releases.
@@ 6. Related Work
The evolution of language models and their integration into existing projects and solutions have opened new horizons in automating routine tasks and processes (Chen et al., 2021; Ziegler et al., 2022). The so-called general-purpose LLMs like Llama 2 (Touvron et al., 2023), Llama 3 (AI@Meta, 2024) or T5 (Raffel et al., 2020) have been adapted to process code-related tasks by fine-tuning on a huge corpus of publicly available source codes and documentation (Code Llama (Rozière et al., 2024) or CodeT5 (Wang et al., 2021)). The latter are pre-equipped with a deep understanding of programming syntax, idioms, and semantic nuances. Such optimized code-specific models can be further fine-tuned on code generation and understanding downstream tasks. Using such a transfer learning approach, these models can reduce the computation time and data quality demands compared to training a new model from scratch while enhancing the model’s ability to generalize across different programs. However, LLMs generally require significantly more computational resources, leading to higher operational costs and energy consumption, thus limiting their application. While our Learner implementation based on a lightweight Transformer model with scaled dot-product attention (Vaswani et al., 2017) mechanism is suitable and sufficient for program behavior modeling and can be trained on a wide range of existing hardware, recent LLMs require access to data-center-grade hardware accelerators. Still, LLMs remain the mainstream of modern software engineering, and, below, we discuss the most relevant applications of neural models in program behavior modeling and the challenges they face.
6.1. Behavior Mocking
@@
There have already been previous attempts to perform behavior mocking by training sequence-to-sequence neural networks. TransCoder (Rozière et al., 2020) represents two sequence-to-sequence models for cross-language translations of code written in C++, Java, and Python with additional support for back translations. Armengol-Estape et al. (Armengol-Estapé and O’Boyle, 2022) train transformer models to replicate the behavior of the GCC compiler. The training data is obtained by selecting C functions from the Anghabench benchmark and compiling them to x86 assembly code. The learned models can correctly compile the C code in less than half of the cases, making replacing compilers with neural machine Translation models invisible in practice. Zaremba et al., in their “Learning To Execute” publication (Zaremba and Sutskever, 2015), try to predict the execution results of short Python programs performing simple arithmetic operations and echo printing by training an LSTM (Hochreiter and Schmidhuber, 1997) model of a Python interpreter on a set of synthesized programs. These works have shown that the current generation of neural network architectures cannot correctly learn the full behavior of such complex programs as compilers or interpreters. While learned models cannot fully replace complex programs, our approach has demonstrated promising results in modeling linear or linearithmic programs. In contrast to previous approaches that rely on sequence-to-sequence models to achieve behavior mocking, Shen. et al. and their Konure (Shen and Rinard, 2021) system relies on the active learning approach to mine the behavior model of applications that interact with relational databases. Konure synthesizes inputs with database content simultaneously, monitors the traffic exchange between the database and the application, and inspects the results of the database queries to continuously refine the application behavior hypothesis. Using the refined application model, Konure can generate the Python code for database querying that is semantically equivalent to the original application code, which might be written in a different programming language. Similar to Konure, our approach learns the program models from the input-output pairs. While Konure is designed purely to model the interaction between an application and a database, our approach does not have such a limitation and can be used to either model the behavior of the whole system or its components and individual functionality.
6.2. Reversing Computation